If we need to bill $509,375 after subtracting a 10% discount, how much does the original total bill need to be?
Answer by Maths68(1474)

Let
Original bill = $x

10% of x = 10/100 (x)
10% of x = 1/10 (x)
10% of x = x/10
Bill = original bill - 10% discount
509375=x-x/10
509375=(10x-x)/10
10* 509375=(10x-x)
5093750=9x
5093750/9=9x/9
565972.20=x
x=565972.20
Original bill = $565972.20